# Upsampling DINOv2 features for unsupervised vision tasks and weakly supervised materials segmentation

### v1

@@ 2 Background
Standard interpolation approaches (bilinear, bicubic, nearest-neighbour) can be used to upsample patch-level features but risks blurring or missing high-resolution details. For ViTs, reducing the stride of the patch embedding layer (creating overlapping patches)16 can increase the feature resolution, but this increases the number of input patches and therefore causes the time and memory scaling problem discussed in Section S5. It can also cause blurring and at very small strides (1 or 2) cause numerical errors.
@@ 3.1 Increasing feature resolution
There are several advantages: this method is model agnostic (assuming they use spatialised features), requires no additional training and works with the strided approach discussed in Section 2. The memory and time cost is linear with the number of transforms $T$ (see Section S5) if batched; for sequential processing the memory cost is constant (relative to the memory and time costs of the strided method). Examples of the feature resolution can be seen in Sections 4.1 and S2. Limitations are discussed in Section S6. Discussion on the impact of these transformations, as well as the number needed for a given stride is available in Section S9.
@@ 3.2 Unsupervised segmentation workflow
These high-resolution features can then be used directly (i.e, without including additional colour information or self-training networks) in unsupervised downstream tasks. Our approach is to produce a ‘Class-Agnostic Segmentation’ (CAS)47 - drawing parallels to class-agnostic object detection - using only the information available in the features, and from that perform tasks like object localization, saliency detection and semantic segmentation. A full diagram of the workflow is shown in Figure 1.
@@
Agglomerative merging with complete linkage is then used to merge clusters that have a distance less than the semantic distance, producing a rough CAS map of the image. This can be refined with a CRF48 to ameliorate the blurring problem. Multiplying the semantic distance by a scalar factor $\lambda$ allows control of the granularity of the merging - varied between 0.95 and 1.1 during experiments. Combining the CAS map and attention information enables various downstream tasks: like object detection by drawing bounding boxes around unconnected regions in each class or saliency detection via attention density of the CAS map.
@@ 3.3 Weakly supervised segmentation
Following existing work in the materials imaging community which train a random forest classifier23, 24, 25, 26 to map from classical image features to user labels, we train a logistic regression classifier to map from our upsampled pixel-wise features from ViT models to user labels. These user labels are from ‘paint-brush’ style annotation, (see Figure 3 for an example) so may be sparse. Each $D$-dimensional vector that describes a (labelled) pixel is used alongside its associated label as a training example for the classifier.
Logisitic regression was chosen for its simplicity and regularizability relative to random forests. We use a random forest classifier for the classical features to ensure similarity to existing schemes 23, 24, 25, 26 and because it showed the best performance.
@@ 4.1 Qualitative comparisons of features
Further comparisons are available in Figure S1. Like FeatUp, our approach works across all models that produce spatialised features, this can be seen in Figure S2. We discuss limitations, including blurring and boundary effects in Figure S7. In each comparison we compare to FeatUp’s fast, single-forward pass JBU approach as it is most similar to the aims of our approach - their implicit approach would produce higher-resolution features but would take far longer.
@@ 4.2 Quantitative feature comparisons via linear probes
To demonstrate the effectiveness of the learned features of DINOv2, a series of linear probes were trained to perform dense tasks like semantic segmentation and depth estimation given bilinear upsamplings of the patch features. Assuming our method produces faithful and useful upsamplings, these linear probes should still be directly applicable to them and produce similar (if not better) results. The results of these linear probes applied to our features for semantic segmentation on VOC201249 is reported in Table S1. There is a performance drop from applying the probe to the features produced by our method, this aligns with FeatUp27’s findings for the strided method of higher resolution but noisier segmentation maps.
@@ 4.3 Unsupervised segmentation experiments
Ours 0.554 0.572
Ours (multi) 0.718 0.725
@@
Ours 0.785 0.654
Table 2: IoU of foreground object segmentation across the CUBS & DUTS datasets. Other values quoted from 21, 18, 52, 17
@@ 4.3.1 Object localization
We then apply this approach to VOC0753 and VOC1249, standard unsupervised object detection benchmarks, reporting the results in Table 1. The success metric is ‘CorLoc’ - the percentage of images where at least one of the predicted boxes has a greater than 50% IoU with at least one of the ground truth boxes. We report two results: the CorLoc when only using the superbox (single object detection) and when using all predicted foreground boxes (multi-object detection). Our method shows comparable performance to state-of-the-art in both single- and multi- object detection.
@@ 4.3.2 Foreground object segmentation
For foreground object segmentation (also called ‘saliency detection’) we use the same approach as in Sections 3.2 and 4.3.1, producing a binary segmentation where all the foreground classes are counted as foreground, the rest as background. We apply our approach to two standard benchmark datasets, CUBS54 and DUTS55, reporting the results in Table 2. Our approach shows good performance on both datasets, surpassing state-of-the-art on DUTS. The success metric is ‘(mean) Intersection over Union’ (mIoU) - the area of the overlap between the prediction for the foreground class and ground truth label, divided by the union of both areas. For multiple classes (i.e, in Section 4.4), this is averaged over the classes.
@@ 4.4 Featuriser for zero-shot weakly supervised segmentation
Figure 3: Explanation of the weakly supervised segmentation workflow, using an example of nucleic segmentation of a TEM image of a Jurkat human T-Cell56. Our method using high-resolution ViT features is able to capture the semantic information (interiority, foreground vs. background) needed for a good segmentation, compared to the classical image features + random forest approach. This disparity in feature richness can be seen in their respective feature spaces.
Following Section 3.3, our aim is to use the richer semantics of these ViT models to perform complex semantic segmentation tasks; as an example we chose a dataset of 135 Transmission Electron Microscopy (TEM) micrographs of human T-cells56 with three classes: background, cell, nucleus. We show an example in Figure 3, where the model using ViT features outperforms the classical method.
@@
Figure 4: Predictions of the trained classifiers using classical or hybrid features on unlabelled examples from the cell dataset. The hybrid features produce good segmentations, demonstrating their ability to generalise well. Note a CRF was used for the hybrid cells and not for the classical.
We note a CRF48 was used to improve the segmentations for the case study, results without a CRF can be found in Section S8. Also included in Section S8 are more example segmentations on unlabelled cells, both with and without a CRF. Our method using the DINOv2 features shows markedly better performance across the dataset than the standard classical features + random forest approach. We attribute this to the richer feature space allowing concepts like interiority and foreground/background distinction to be expressed.
@@
The first micrograph is from a high resolution Scanning Electron Microscope (SEM) image of lithium nickel manganese cobalt oxide cathode57, 58 - the ViT features are able to distinguish between in-plane (light grey, flat) and out-of-plane (light grey, textured) active material (AM) present due to the ‘pore-back’ effect57, 58. The ViT features also admit the distinction between voids inside an AM particle and the out-of-plane AM (which have similar textures).
@@
We note that these experiments were performed using the default Weka feature set23; choosing a set more suited to the problem may improve performance, though this represents significant trial-and-error. Discussion of the the limitations can be found in Section S6, focusing on the practical implications of using these features in a materials context.
Figure 5: Application of classical, DINOv2 and classical + DINOv2 (hybrid) features for weakly supervised segmentation of materials such as NMC cathodes57, 58, SiC oxide layers59 and cast iron alloys60. The hybrid scheme is able to combine the strong semantic features of DINOv2 with the high-resolution classical features for good segmentations of complex tertiary and quaternary phases.
@@ 5 Conclusion
To conclude, we have demonstrated a novel single-forward pass upsampling technique for features of vision transformers, like DINOv2. This was then applied to perform unsupervised object detection and segmentation, achieving strong results compared to baselines. One potential use for this unsupervised workflow is generating high-quality semantic region proposals for interactive segmentation datasets 10, 61.
@@ Acknowledgements
Code Availability
The code needed to reproduce the results of the paper is available at https://github.com/tldr-group/HR-Dv2 with an MIT license agreement.
@@
S1 Linear probe evaluation
Method Pascal VOC
Bilinear 0.806
FeatUp (JBU) 0.825
Ours 0.701
Table S1: Drop-in performance when applying a linear classifier trained on bilinear-upsampled DINOv2-S-14 features to FeatUp upsampled features and features upsampled using our method (strided + shift transforms).
@@
S3 Averaging positional effects
Figure S3: Demonstrating the positional effects present when striding DINOv2 features in (a) a Scanning Electron Microscope (SEM) image of a solid-oxide fuel cell and (b) a photo of a dog. These can be ameliorated by including flip transformations during the upsampling process, which can be used in conjunction with shift transforms at the cost of computing $N_{\text{shift}}\times N_{\text{flip}}$ feature maps, allowing the features to express more semantic information when displayed in a 3-component PCA. The input image size was (518, 518). Note there is still a center-image bias for the SOFC even after flip transforms have been applied.
S4 Small object retrieval and keypoint matching
Figure S4: Small object retrieval (red cross) and keypoint matching (coloured dots) for various feature upsampling methods with DINOv2-S-14, stride 7. Following FeatUp, we compute high-resolution features for a query and target image, then find the most similar point in the target image to a given query point (and additional keypoints). The colourmap indicates similarity to the query point, red for similiarity, blue for dissimilarity. We note our similarities are more localised than FeatUp’s (JBU).
@@
S5 Practical considerations
Figure S6: Memory and time usage of various upsampling methods as function of square image length for DINOv2-S-14. We fix the stride for both strided and our methods to be 4 and use shift transforms with a distance of 2 and in an 8-neighbourhood for a total of 17 transformations. ‘Batch’ refers to computing the features of the transformed image batch in one forward pass, ‘Sequential’ to computing them one at a time. We note JBU has a harsher memory scaling as image length increases but far better time scaling. We achieve this better memory scaling than the expected $n^{2}$ via memory efficient attention62, which can be added post-hoc into any existing ViT network and admits $\sqrt{n}$ memory cost with $n$ tokens (though still an $n^{2}$ time cost). If the stride were to be increased beyond 4, the memory cost would exceed JBU, though this is undesirable due to numerical errors encountered at low strides/high upsampling factors. Float16 precision was used for both measurements; note for image length $l$, $n\propto l^{2}$. Values measured on an NVIDIA RTX A6000.
S6 Limitations
Figure S7: We highlight the limitations of our feature upsampling with two example images, the most major being the blurring caused by reducing the stride of the patch embedding layer - this is most evident around the plant pot’s edges when compared to FeatUp’s JBU. Comparing our approach applied to DINO-S-8 vs DINOv2-S-14 shows that the resolution is best improved by using models with a smaller patch size, which relies on the group training the feature model; FeatUp does not suffer this problem. One minor problem is edge effects caused by using pixel-shifts with wrap boundary conditions - this causes information from opposite sides of the image to mix, albeit for a small ($<4$) number of pixels. Another problem is the increasing time-cost as a function of image resolution (see Figure S6), which can prove prohibitive for real-time applications in domains with high-resolution images, like materials science.
The primary limitation of our feature upsampling method is the blurring introduced by the striding (see Figure S7), which limits the resolution of the later downstream tasks. There are also practical problems with decreasing the model stride (and therefore increasing the number of tokens, $n$): although the memory cost does not scale $n^{2}$ thanks to memory-efficient attention62, the time-cost still scales with $n^{2}$ (i.e, $N_{\text{pixels}}^{4}$). See Figure S6 for more discussion. Materials science frequently handles high resolution volumetric data - reducing the featurisation time is important to ensure these data can be processed.
The unsupervised segmentation workflow tends to decompose large foreground objects into parts (head/body, etc.) and treat those as separate classes. This could be improved by merging the features of individual pixels (rather than clusters), but this would be expensive computationally.
The resolution problems and time-cost of generating the high-resolution features limit the applicability of the method in user-facing weakly supervised materials segmentation. Furthermore, it is possible for the classifiers to overfit to positional information present in the deep features - this requires additional homogeneous labelling to overcome or some augmentations to reduce their importance, see Supplementary Section S3. Empirically, combining classical and deep features can make the trained classifiers ‘brittle’, i.e, less sensitive to semantic information and therefore worse at generalizing.
@@
S7 Unsupervised segmentation examples
Figure S9: Examples of the unsupervised segmentation workflow applied to a natural scale image and an SEM of mesoporous KIT-6 silica63, showing how the unsupervised semantic map and attention densities can be used for object detection and localization.

### v2

@@ 2 Background
DINOv213 iterated on the DINO formula, including iBOT loss and various training stability improvements. It trained to learn features over a curated dataset of around 140M images and showed strong performance across various benchmarks. Evaluation was done by freezing the model and training small head networks (linear probes for classification or convolutional layers for dense tasks) on the features.
@@
Standard interpolation approaches (bilinear, bicubic, nearest-neighbour) can be used to upsample patch-level features but risks blurring or missing high-resolution details. For ViTs, reducing the stride of the patch embedding layer (creating overlapping patches)16 can increase the feature resolution, but this increases the number of input patches and therefore causes the time and memory scaling problem discussed in Section 5. It can also cause blurring and at very small strides (1 or 2) cause numerical errors.
@@ 3.1 Increasing feature resolution
There are several advantages: this method is model agnostic (assuming they use spatialised features), requires no additional training and works with the strided approach discussed in Section 2. The memory and time cost is linear with the number of transforms $T$ (see Section S7.2) if batched; for sequential processing the memory cost is constant (relative to the memory and time costs of the strided method). Examples of the feature resolution can be seen in Sections 4.1 and S2. Limitations are discussed in Section 5. Discussion on the impact of these transformations, as well as the number needed for a given stride is available in Section S7.
In terms of the Vision Transformer whose features we upsample, we choose either DINO-S-8 or DINOv2-S-14. DINO-S-8 has a smaller patch size, and therefore blurs less when upsampled, but was trained on a smaller dataset and has less strong semantic distinctions than DINOv2. Other ViTs were tested (see Supplementary Figure S2) but were fixed in their input size. We used the DINOv2 checkpoint with added register tokens, which prevents some background tokens from having anomalously high attention values 47.
@@ 3.2 Unsupervised segmentation workflow
These high-resolution features can then be used directly (i.e, without including additional colour information or self-training networks) in unsupervised downstream tasks. Our approach is to produce a ‘Class-Agnostic Segmentation’ (CAS)48 - drawing parallels to class-agnostic object detection 17 - using only the information available in the features, and from that perform tasks like object localization, saliency detection and semantic segmentation. A full diagram of the workflow is shown in Figure 1.
@@
Agglomerative merging with complete linkage is then used to merge clusters that have a distance less than the semantic distance, producing a rough CAS map of the image. This can be refined with a CRF49 to ameliorate the blurring problem. Multiplying the semantic distance by a scalar factor $\lambda$ allows control of the granularity of the merging - varied between 0.95 (less merging) and 1.1 (more merging) during experiments. Increasing the number of initial clusters increases the time and memory complexity of the merging step, decreasing reduces the object-level resolution and makes the statistics of the cluster ‘attention density’ less robust; we choose 80 as a middle ground.
Combining the CAS map and attention information enables various downstream tasks: like object detection by drawing bounding boxes around unconnected regions in each class or saliency detection via attention density of the CAS map. For each unsupervised segmentation experiment in Sections 4.3.1 and 4.3.2 we used DINOv2-S-14 with stride set to 4, shifts of 1 and 2px and flip transforms. The merge parameter $\lambda$ was set to 1.0. Stride 4 was the highest resolution achievable accounting for memory and time scaling, shifts greater than 2 had little affect on the feature resolution (see Section S7) and flip transformations averaged out some positional bias (see Figure 9).
@@ 3.3 Weakly supervised segmentation
Following existing work in the materials imaging community which train a random forest classifier23, 24, 25, 26 to map from classical image features to user labels, we train a classifier to map from our upsampled pixel-wise features from ViT models to user labels. These user labels are from ‘paint-brush’ style annotation, (see Figure 4 for an example) so may be sparse. Each $D$-dimensional vector that describes a (labelled) pixel is used alongside its associated label as a training example for the classifier.
We compare across three ‘feature-sets’: Weka style classical features (Gaussian blurs, Sobel filter, Hessian filter, membrane projections, difference of Gaussians), upsampled DINOv2-S-14 with stride 4 and a ‘hybrid’ scheme, where we concatenate the classical and upsampled ViT features in the channel dimension. This ‘hybrid’ approach allows the classifier to leverage both the high-resolution (but inexpressive) classical features and the coarser, but more semantic upsampled ViT features. This hybrid approach is therefore a superset of the classical features, so should have strictly improved performance (assuming the upsampled ViT features add useful information).
Across the experiments in Section 4.4 we focus on two types of classifier: random forests and logistic regression. We use a random forest classifier for the classical features to ensure similarity to existing schemes 23, 24, 25, 26. For the upsampled ViT features we also experiment with using logistic regression, chosen for its simplicity relative to random forests, with the goal of demonstrating that the more expressive (and non-linear) features required a less complex classifier. We justify this choice in Figure S12, which shows the segmentation performance when using upsampled ViT features to be fairly consistent across classifiers, and for the classical features to perform best when using random forests.
Unless otherwise stated, the random forest is ‘Weka-style’, with 200 trees, 2 features per split and depth of 10. The logistic regression used default scikit-learn 50 parameters, with L2 regularization and 1000 iterations. Both classifiers had class-frequency weights applied. For ViT upsampling, we used DINOv2-S-14 with stride 4, shifts of 1 and 2px and flip transforms.
3.4 Datasets
We test our unsupervised segmentation workflow on two tasks: object localization - determining a bounding box for every foreground object in an image - and foreground object segmentation, which is predicting all pixels that belong to the foreground object(s) (i.e, a binary segmentation). For the object localization task (Section 4.3.1) we use the Video Object Classification (VOC) datasets, specifically VOC07 and VOC1251, 52. These are standard natural images of a range of everyday objects, where each image contains one or more examples of an object and the corresponding bounding boxes. It is also possible to use VOC as a semantic segmentation dataset, which we do to evaluate the performance of the upsampled features with linear probes in Section 4.2.
For the foreground object segmentation (Section 4.3.2) we use two datasets: the Caltech-UCSD Birds dataset (CUBS)53 and the Dalian University of Technology (DUTS) dataset54. CUBS contains images of various birds in difference contexts, alongside binary segmentations of the bird vs the background. DUTS contains images of various everyday objects and their segmentation.
When evaluating the performance in weakly-supervised segmentation in Section 4.4 we use a hand-labelled dataset of Transmission Electron Micrographs (TEM) of human T-cells 55 with three classes: background, cell, nucleus. We also examine a series of micrographs from materials science taken with a variety of instruments. These images do not have a ground truth.
3.5 Evaluation Metrics
Segmentation performance was evaluated using Intersection over Union (IoU), which for a binary segmentation (foreground vs background) measures the overlap of the predicted foreground pixels and ground truth label, divided by the total area of the two predictions. This ranges between 0 (no overlap) and 1 (predicted and ground truth are the same). For multiple classes, the IoU can be calculated separately for each class (by treating all other classes as background) and averaged, producing the mean Intersection over Union, or mIoU.
We measure object detection performance using ‘Correct Localization’ or CorLoc, which is ‘true’ if at least one predicted bounding boxes for an image has a greater than 50% IoU with at least of the ground truth bounding boxes. This is averaged over all images in the dataset to get a score, which is the fraction of all images with a correct localization.
@@ 4.1 Qualitative comparisons of features
Further comparisons are available in Figure S1. Like FeatUp, our approach works across all models that produce spatialised features, this can be seen in Figure S2. We discuss limitations, including blurring and boundary effects in Figure 7. In each comparison we compare to FeatUp’s fast, single-forward pass JBU approach as it is most similar to the aims of our approach - their implicit approach would produce higher-resolution features but would take far longer.
@@ 4.2 Quantitative feature comparisons via linear probes
Method VOC12
Bilinear 0.806
FeatUp (JBU) 0.825
Ours (DINOv2-S-14) 0.701
Table 1: mIoU on the VOC12 dataset when applying a linear classifier trained on bilinear-upsampled DINOv2-S-14 features to FeatUp upsampled features and features upsampled using our method (strided + shift transforms).
To demonstrate the effectiveness of the learned features of DINOv2, a series of linear probes were trained to perform dense tasks like semantic segmentation and depth estimation given bilinear upsamplings of the patch features 13. Assuming our method produces faithful and useful upsamplings, these linear probes should still be directly applicable to them and produce similar (if not better) results. The results of these linear probes applied to our features for semantic segmentation on VOC201252 is reported in Table 1. There is a performance drop from applying the probe to the features produced by our method, this aligns with FeatUp27’s findings for the strided method of higher resolution but noisier segmentation maps.
@@ 4.3 Unsupervised segmentation experiments
Figure 3: Examples of the unsupervised segmentation workflow applied to a natural scale image and an SEM of mesoporous KIT-6 silica56, showing how the unsupervised semantic map and attention densities can be used for object detection and localization.
@@
Ours (DINOv2-S-14) 0.554 0.572
Ours (DINOv2-S-14, multi) 0.718 0.725
@@
Ours (DINOv2-S-14) 0.785 0.654
Table 3: IoU of foreground object segmentation across the CUBS & DUTS datasets. Other values quoted from 21, 18, 59, 17. If a paper does not report a value for a dataset it has been replaced with a hyphen.
@@ 4.3.1 Object localization
We then apply this approach to VOC0751 and VOC1252, standard unsupervised object detection benchmarks, reporting the results in Table 2. The success metric is ‘CorLoc’ - the percentage of images where at least one of the predicted boxes has a greater than 50% IoU with at least one of the ground truth boxes. We report two results: the CorLoc when only using the superbox (single object detection) and when using all predicted foreground boxes (multi-object detection). Our method shows comparable performance to state-of-the-art in both single- and multi- unsupervised object detection.
@@ 4.3.2 Foreground object segmentation
For foreground object segmentation (also called ‘saliency detection’) we use the same approach as in Sections 3.2 and 4.3.1, producing a binary segmentation where all the foreground classes are counted as foreground, the rest as background. We apply our approach to two standard benchmark datasets, CUBS53 and DUTS54, reporting the results in Table 3. Our approach shows good performance on both datasets, surpassing state-of-the-art on DUTS.
@@ 4.4 Featuriser for zero-shot weakly supervised segmentation
Figure 4: Explanation of the weakly supervised segmentation workflow, using an example of nucleic segmentation of a TEM image of a Jurkat human T-Cell55. Our method using high-resolution ViT features is able to capture the semantic information (interiority, foreground vs. background) needed for a good segmentation, compared to the classical image features + random forest approach. This disparity in feature richness can be seen in their respective feature spaces.
Application to cell TEM dataset
Following Section 3.3, our aim is to use the richer semantics of these ViT models to perform complex semantic segmentation tasks; as an example we chose a dataset of 135 Transmission Electron Microscopy (TEM) micrographs of human T-cells55 with three classes: background, cell, nucleus. We show an example in Figure 4, where the model using ViT features outperforms the classical method.
@@
We note a CRF49 was used to improve the segmentations for the case study, results without a CRF can be found in Section S6. Also included in Section S6 are more example segmentations on unlabelled cells, both with and without a CRF. Our method using the DINOv2 features shows markedly better performance across the dataset than the standard classical features + random forest approach. We attribute this to the richer feature space allowing concepts like interiority and foreground/background distinction to be expressed. A figure showing the resolution increase as the stride is reduced and transforms are added is available in Section S7.1.
Figure 5: Predictions of the trained classifiers using classical or hybrid features on unlabelled examples from the T-cell dataset. The hybrid features produce good segmentations, demonstrating their ability to generalise well. Note a CRF was used for the hybrid cells and not for the classical.
Results on diverse materials micrographs
Figure 6: Application of classical, DINOv2 and classical + DINOv2 (hybrid) features for weakly supervised segmentation of materials such as NMC cathodes60, 61, SiC oxide layers62, cast iron alloys63, glutamic acid polymorphs 64, a nickel superalloy 62 and gold nanoparticles 65. The hybrid scheme is able to combine the strong semantic features of DINOv2 with the high-resolution classical features for good segmentations of complex tertiary and quaternary phases.
@@
The first micrograph is from a high resolution Scanning Electron Microscope (SEM) image of lithium nickel manganese cobalt oxide cathode60, 61 - the ViT features are able to distinguish between in-plane (light grey, flat) and out-of-plane (light grey, textured) active material (AM) present due to the ‘pore-back’ effect60, 61. The ViT features also admit the distinction between voids inside an AM particle and the out-of-plane AM (which have similar textures).
@@
The fourth micrograph is a slice from an X-ray Computed Tomography (XCT) of glutamic acid64, displaying two crystal polymorphs: the larger, blocky $\alpha$-polymorph and the smaller, needle-like $\beta$-polymorph64. The classical features are unable to distinguish between the two phases, struggling to delineate between edges of a large $\alpha$ block and the $\beta$ needles, whilst the DINOv2 features succeed but are blurred. The combination of the two produces a good segmentation of both phases - this ability to classify using structural information (rather than focussing on pixel-value) is useful in distinguishing between polymorphs for characterization techniques that have elemental contrast (like XCT or SEM), as the polymorphs will have similar pixel-values.
The fifth micrograph is an SEM image of a nickel-based superalloy62 with three phases: a large, connected matrix phase, secondary precipitates (large blobs) and tertiary precipitates (small blobs). Again, we find that the hybrid approach produces a better segmentation, especially of the third phase. The final micrograph is a high-resolution TEM image of gold nanoparticles 65, which displays noise and an exposure variation across the image that the classical features struggle to capture (given the localised labels).
@@
We note that these experiments were performed using the default Weka feature set23; choosing a set more suited to the problem may improve performance, though this represents significant trial-and-error. A random forest classifier was used for all segmentations, and the labels were kept the same i.e, the only difference between the columns is the feature-set used. Labels were produced using the SAMBA web-app66. Discussion of the limitations of the deep features can be found in Section 5, focusing on the practical implications of using them in a materials context.
5 Limitations
Figure 7: We highlight some of the limitations of our upsampling, primarily the blurring caused by reducing the model stride. This is reduced for models with a smaller base patch size (DINO-S-8 vs DINOv2-S-14), though a smaller base patch size incurs a higher memory cost; FeatUp (JBU) suffers some blurring, but retains sharp boundaries for foreground objects.
Figure 8: Memory and time usage of various upsampling methods as function of square image length for DINOv2-S-14. We fix the stride for both strided and our methods to be 4 and use shift transforms with a distance of 2 and in an 8-neighbourhood for a total of 17 transformations. ‘Batch’ refers to computing the features of the transformed image batch in one forward pass, ‘Sequential’ to computing them one at a time. We note JBU has a harsher memory scaling as image length increases but far better time scaling. We achieve this better memory scaling than the expected $n^{2}$ via memory efficient attention67, which can be added post-hoc into any existing ViT network and admits $\sqrt{n}$ memory cost with $n$ tokens (though still an $n^{2}$ time cost). If the stride were to be increased beyond 4, the memory cost may exceed JBU, though this is undesirable due to numerical errors encountered at low strides/high upsampling factors. Float16 precision was used for both measurements; note for image length $l$, $n\propto l^{2}$. Values measured on an NVIDIA RTX A6000.
The primary limitation of our feature upsampling method is the blurring introduced by the striding (see Figure 7), which limits the resolution of the later downstream tasks. There are also practical problems with decreasing the model stride (and therefore increasing the number of tokens, $n$): although the memory cost does not scale $n^{2}$ thanks to memory-efficient attention67, the time-cost still scales with $n^{2}$ (i.e, $N_{\text{pixels}}^{4}$).
We present a comparison of time and memory usage of the various upsampling techniques as function of image side length in Figure 8. Our approach has a better memory scaling than JBU, but a worse time scaling; in an application context where memory is fixed, a higher time cost is preferable. However, materials science frequently handles high resolution volumetric data, so reducing this featurisation time is important to ensure these data can be processed.
The unsupervised segmentation workflow tends to decompose large foreground objects into parts (head/body, etc.) and treat those as separate classes - see Figure S5. This could be improved by merging the features of individual pixels (rather than clusters), but this would be expensive computationally.
The resolution problems and time-cost of generating the high-resolution features limit the applicability of the method in user-facing weakly supervised materials segmentation. Furthermore, it is possible for the classifiers to overfit to positional information present in the deep features - this requires additional homogeneous labelling to overcome, or some augmentations to reduce their importance, shown in Figure 9. Empirically, combining classical and deep features can make the trained classifiers ‘brittle’, i.e, less sensitive to semantic information and therefore worse at generalizing.
Figure 9: Demonstrating the positional biases present in DINOv2 features, and their consequences on weakly-supervised segmentation. The deep features, in contrast to the classical features, contain positional information, which the classifier can fit to (in the case of non-homogenous labels). Adding flip transformations can ameliorate this, though note there is still a center-image bias even after flip transforms have been applied.
@@ 6 Conclusion
To conclude, we have demonstrated a novel single-forward pass upsampling technique for features of vision transformers, like DINOv2. This was then applied to perform unsupervised object detection and segmentation, achieving strong results compared to baselines. One potential use for this unsupervised workflow is generating high-quality semantic region proposals for interactive segmentation datasets 10, 68.
@@ Acknowledgements
Code and Data Availability
The code needed to reproduce the results of the paper is available at https://github.com/tldr-group/HR-Dv2 with an MIT license agreement. The data for the case studies is available at https://zenodo.org/records/14722134.
@@
S1 Full algorithms
1:Image II (of size $(C,H,W)$), a ViT feature extraction model G with hidden dimension $k$ and number of attention heads $N_{h}$, a set of image transformations (pixel shifts, flips) as partial functions $T$ with $N_{T}=|T|$, which last-layer attention vector (if any) to upsample $(\textbf{q}\text{uery},\textbf{k}\text{ey},\textbf{v}\text{alue or }\textbf{o}\text{utput})$. Note this is the attention map for the global [CLS] token.
2:Reduce stride of G’s input patch projection layer from 16 or 14 down to some $S$ (usually $S=4$).
3:
[…]
8:$F$, shape $(k,H,W)$ & AA, shape $(N_{h},H,W)$ $\leftarrow$ Average over all $f\in F^{\prime}$ up to $k^{th}$ channel for features, and average all $f\in F^{\prime}$ from $k^{th}$ to $k+N_{h}^{th}$ channel for attention. return $F$, the upsampled ViT features of image $I$, AA the upsampled attention maps of II.
Algorithm 1 Transformation based ViT feature upsampling (batched)
1:Image II (of size $(C,H,W)$), number of k-means clusters $C$, associated upsampled features $F$ (shape $(k,H,W)$) and attention maps AA (shape $(N_{h},H,W)$) of the [CLS] token. Optional number of classes $N_{c}$ and merge threshold multiplier $\lambda$.
2:Sum the attention maps across the heads to produce $A^{\prime}$, shape $(H,W)$
3:
[…]
9:if $N_{c}$ supplied then
10:  $S_{M}$ $\leftarrow$ Greedily merge clusters in $\mu$ based on cosine distance until $N_{c}$ clusters remain
11:else
12:  $S_{M}$ $\leftarrow$ Greedily merge clusters in $\mu$ if their cosine distance is less than the semantic distance i.e, $\text{dist}(F_{\mu_{i}},F_{\mu_{j}})$ < $d_{s}$ to get segmentation
[…]
Algorithm 2 Unsupervised ViT feature-clustering based segmentation
1:Image II (of size $(C,H,W)$), sparse user label mask $L$ of size $(H,W)$, desired featureset FSFS (one of ‘Classical’, ‘Upsampled ViT’, ‘Hybrid’), a classifier $f$ - can be linear, logistic, random forest, XGB, etc.
2:
3:if $FS=\text{`Classical'}$ then
4:  $F$, shape $(N_{c},H,W)$ $\leftarrow$ Classical Weka-style features (edges, Gaussian blurs, …)
5:else if $FS=\text{`Upsampled ViT'}$ then
6:  $F$, shape $(k,H,W)$ $\leftarrow$ Get upsampled ViT features from Algorithm 1
7:else if $FS=\text{`Hybrid'}$ then
8:  $F_{c}$, shape $(N_{c},H,W)$ $\leftarrow$ Classical Weka-style features (edges, Gaussian blurs, …)
9:  $F_{v}$, shape $(k,H,W)$ $\leftarrow$ Get upsampled ViT features from Algorithm 1
10:  $F$, shape $(N_{c}+k,H,W)$ $\leftarrow$ concat $F_{c}$ and $F_{v}$ in the channel dimension
11:$(X,Y)$ $\leftarrow$ get features of all labelled pixels $X$ from $F$ and their class values $Y$ from $L$; optionally sample; shuffle
12:$f$ $\leftarrow$ fit classifier $f$ over training data $(X,Y)$
13:$S$, shape $(H,W)$ $\leftarrow$ apply trained classifier $f$ to features of unlabelled pixels, predicting a class for each one return $S$, segmentation map of II given user class labels in $L$ and desired featureset FSFS and $f$, the trained classifier for re-use.
Algorithm 3 Weakly supervised segmentation with upsampled ViT features
@@
Figure S1 shows a comparison of feature maps across different upsampling methods for a set of natural images.
Figure S2 shows a comparison of feature maps across different ViT backbones using our upsampling method (stride 4, shifts 1,2, flips). We see that models with a smaller patch size (DINO-S-8) achieve higher resolution, and that DINOv2-S-14 suffers the worst blurring, moreso than models with a larger patch size (DINO-S-16, DEiT-s 16, ViT-s 16).
@@
S3 Small object retrieval and keypoint matching
Following FeatUp 27, we perform keypoint matching and similarity visualisation using the upsampled features in Figures S3 and S4. In it, the feature vectors of series of keypoints on a query image (red cross, coloured dots) are taken, and the most similar points (in terms of cosine similarity) are found in a target image. We also plot the cosine similarity of every pixel to the feature vector red cross as a heatmap, red for similar, blue for dissimilar. This gives a measure of how localised the similarities are across the feature map in the target image. We see that the similarities are more localised with our method than Featup (JBU).
Figure S3: Small object retrieval (red cross) with additional keypoint matching (blue, orange, green & red coloured dots) for various feature upsampling methods with DINOv2-S-14, stride 7. Following FeatUp, we compute high-resolution features for a query (left, stop sign) and target image (right), then find the most similar point in the target image to a given query point (red cross) and additional keypoints). The colourmap indicates similarity to the query point, red for similarity, blue for dissimilarity. The coloured dots are simply additional keypoints applied to the query image - we look for the most similar vector in the target image across the different upsampling methods. We note our similarities are more localised than FeatUp’s (JBU).
@@
S4 Failure modes of the unsupervised workflow
We present failure modes of the unsupervised workflow in Figure S5. The first is overdecomposing objects (i.e the dog’s body and leges), the second is merging an object with its shadow and the last two are examples of over-merging (the horse/bike and its rider). These are, to some degree, functions of the relativity of the DINO features and the positional biases. Objects close in space to one another will be similar in terms of their positional information and foregroundedness, so their cosine distance will be smaller (and therefore be more likely to be merged).
@@
S5 Unsupervised segmentation examples
We show some examples of the unsupervised segmentation in Figure S6, which demonstrates both the semantic segmentation from the determined semantic distance, and bounding boxes around the foreground classes based on the the attention density. The approach is able to separate touching objects of different classes (eagle, branch) and merge separate instances of the same class (couch, leaves, bears). These examples were produced using DINOv2-S-14, stride 4, shifts of 1 and 2px and flip transforms.